Solve.
1) square root of 6x-3 EQUALS 3
2) square root of 3x-8 PLUS 4 = 6

Answer by edjones(8007)

sqrt(6x-3)=3
sqrt(6x-3)^2=3^2
6x-3=9
6x-3+3=9+3
6x=12
6x/6=12/6
x=2
.
Ed